OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. Show all work for full credit. 1, -4, -9, -14, . . .

OpenStudy (amistre64):

define a rule of expression that defines the sequence

OpenStudy (amistre64):

once you know how to determine {a_n} then summing is simple enough

OpenStudy (anonymous):

\[S _{n}=\frac{ n }{ 2 } [2a + (n-1)d]\]

OpenStudy (amistre64):

for an arithmetic sequence:\[\sum_{n=a}^b(a_1+d(n-1))=\frac{b(a_1+a_b)}{2}\]

OpenStudy (amistre64):

my choice of letters wasnt good lol

OpenStudy (anonymous):

n=is the number of terms a=1st term d=common difference Sn=Sum of the n terms

OpenStudy (amistre64):

\[\sum_{n=p}^q(a_p+d(n-1))=\frac{q(a_p+a_q)}{2}\] might be better

OpenStudy (amistre64):

ugh ... the range is not q, but q-p ... oh well